# Nonhermitian topological zero modes at smooth domain walls: Exact solutions

### v1

@@ Abstract
The bulk-boundary correspondence predicts the existence of boundary modes localized at the edges of topologically nontrivial systems. The wavefunctions of hermitian boundary modes can be obtained as the eigenmode of a modified Jackiw-Rebbi equation. Recently, the bulk-boundary correspondence has been extended to nonhermitian systems, which describe physical phenomena such as gain and loss in open and non-equilibrium systems. Nonhermitian energy spectra can be complex-valued and exhibit point gaps or line gaps in the complex plane, whether the gaps can be continuously deformed into points or lines, respectively. Specifically, line-gapped nonhermitian systems can be continuously deformed into hermitian gapped spectra. Here, we find the analytical form of the wavefunctions of nonhermitian boundary modes with zero energy localized at smooth domain boundaries between topologically distinct phases, by solving the generalized Jackiw-Rebbi equation in the nonhermitian regime. Moreover, we unveil a universal relation between the scalar fields and the decay rate and oscillation wavelength of the boundary modes. This relation quantifies the bulk-boundary correspondence in nonhermitian line-gapped systems in terms of experimentally measurable physical quantities and is not affected by the details of the spatial dependence of the scalar fields. These findings shed some new light on the localization properties of boundary modes in nonhermitian and topologically nontrivial states of matter.
@@ I Introduction
The bulk-boundary correspondence, which predicts the presence and number of localized boundary modes in relation to the topological invariants in topologically nontrivial condensed matter systems, such as topological insulators and superconductors schnyder_classification_2009 ; hasan_colloquium:_2010 ; qi_topological_2011 ; shen_topological_2011 ; shen_topological_2017 , regardless of the specific details of the physical laws governing the system hatsugai_chern_1993 ; ryu_topological_2002 ; teo_topological_2010 . However, only the specific form of these laws can accurately describe the local physical properties of these excitations, such as their localization and spatial behavior. Generalizations of the bulk-boundary correspondence apply to nonhermitian systems ashida_non-hermitian_2020 ; okuma_non-hermitian_2023 as well. Nonhermitian Hamiltonians naturally describe a variety of physical systems, such as open systems in the presence of gain and loss mechanisms, or dissipation, and exhibit the presence of complex energy spectra shen_topological_2018 ; gong_topological_2018 ; kawabata_symmetry_2019 ; ashida_non-hermitian_2020 ; okuma_non-hermitian_2023 ; schindler_hermitian_2023 ; nakamura_bulk-boundary_2024 in several different condensed matters and metamaterials gao_observation_2015 ; ghatak_observation_2020 ; zhang_observation_2021 ; su_direct_2021 ; fleckenstein_non-hermitian_2022 ; lin_observation_2022 ; liu_experimental_2023 ; wang_non-hermitian_2023 ; ochkan_non-hermitian_2024 . A fundamental distinctions in this case is between spectra with line gaps and point gaps shen_topological_2018 ; gong_topological_2018 ; kawabata_symmetry_2019 . Point gaps describe energy eigenvalues that do not cross a reference point (usually the origin at zero), while line gaps describe energy eigenvalues that do not cross a reference line in the complex plane. Any nonhermitian Hamiltonian with line gaps can be adiabatically deformed into a corresponding hermitian hamiltonian with conventional energy gaps. Hence, the bulk-boundary correspondence and the definition of topological invariants can be directly extended in this case. Hermitian Hamiltonians with nontrivial topology are naturally described by a modified Jackiw-Rebbi equation shen_topological_2011 ; shen_topological_2017 , which is a spinorial second-order linear differential equation, which generalized the original Jackiw-Rebbi equation jackiw_solitons_1976 , which is a spinorial first-order linear differential equation. Henceforth, nonhermitian hamiltonians can be described by a nonhermitian version of the modified Jackiw-Rebbi equation with complex coefficients.
@@
In recent works, we derived the analytical solutions of the modified Jackiw-Rebbi equation describing zero energy boundary modes at smooth domain walls between topological trivial and nontrivial phases in the hermitian regime marra_topological_2024 and extended some of these findings to the nonhermitian regime marra_zero_2025 . In this work, we derive the analytical form of the wavefunctions of zero modes in nonhermitian line-gapped systems described by a modified Jackiw-Rebbi equation with space-dependent mass and velocity terms, allowing for complex coefficients. The space dependency describes the case of a smooth domain wall, i.e., the case where the nonuniform space dependence is confined in a region of finite length $w$. This allows us to derive the bulk-boundary correspondence for nonhermitian line-gapped systems by directly establishing the correspondence between the nonhermitian topological invariant and the asymptotic decay rate of the zero modes. We thus analyze several properties of these zero modes, such as the localization length $\xi$, oscillation wavelength $\lambda$, conditions for the reality of the wavefunction, local properties, and local topological invariant, and show specific examples of nonhermitian zero modes corresponding to different regimes. This allows us to distinguish between qualitatively different cases: featureless zero modes in the limit $w\to 0$, having "no hair" (in analogy with other featureless objects such as black holes); non-featureless zero modes in the regime $\xi,\lambda>w>0$, having "short hair" (i.e., effectively featureless at long length scales); non-featureless zero modes in the regime $w>\xi,\lambda>0$ having "long hair" (i.e., being nonfeatureless at all length scales). We also discuss possible experimental signatures of the bulk-boundary correspondence by unveiling a universal relation between the local properties of the zero modes (localization length and oscillations wavelength $\lambda$) and bulk physical quantities.
II Zero modes: Exaxt solutions
II.1 The modified Jackiw-Rebbi equation in the nonhermitian regime
@@
Figure 1: Energy spectra of the generalized Jackiw-Rebbi equation with uniform fields $v(x)=v$, $m(x)=m$ in Eq. 4. (a) Hermitian case with $v,m\in\mathbb{R}$ corresponding to a real energy spectrum with an energy gap. (b) nonhermitian case with $v,m\notin\mathbb{R}$ corresponding to a complex energy spectrum with a line gap.
@@
The Jackiw-Rebbi equation jackiw_solitons_1976 is a Dirac equation where the mass term is space-dependent. Here, we consider a generalization of the Jackiw-Rebbi equation, where both the mass and Dirac velocity are space-dependent, and with an additional quadratic term in the momentum given by
@@
where $\Psi(x)^{\dagger}=[\psi(x)^{\dagger},\psi(x)]$ is the fermionic field, $\sigma_{xyz}$ the Pauli matrices, $p=-\mathrm{i}\partial$ the momentum operator, $\eta\geq 0$. Furthermore, we consider here the nonhermitian case where the fields are complex-valued ${m}(x),{v}(x)\in\mathbb{C}$. This equation can be interchangeably referred to as a modified Jackiw-Rebbi equation or a modified Dirac equation with space-dependent mass and Dirac velocity. The hermitian case ${m}(x),{v}(x)\in\mathbb{R}$ describes topological insulators and superconductors and has been studied at length in our previous work Ref. marra_topological_2024, . Here, we will address the cases where $\mathrm{Im}({m}(x))\neq 0$ or $\mathrm{Im}({v}(x))\neq 0$, which generalize the equation above to the nonhermitian case. The nonhermitian case $\mathrm{Im}({m}(x))\neq 0$ may describe, e.g., a topological insulator or superconductor subject to a nonhermitian potential $V(x)={m}(x)$, while $\mathrm{Im}({v}(x))\neq 0$ may describe, e.g., a topological superconductor with a nonhermitian superconducting order parameter describing dissipation effects on Cooper pairs. We will consider the zero modes $E=0$ on infinite or semi-infinite intervals that are normalizable and satisfy Dirichlet boundary conditions. The zero modes of the generalized Jackiw-Rebbi equation in Eq. 1 are eigenstates of $\sigma_{x}$, i.e.,
$\Psi(x)\propto\begin{pmatrix}1\\ s\end{pmatrix}\varphi(x),\,\,\sigma_{x}\Psi(x)=s\Psi(x),$ (2)
for $s=\pm 1$ and with $\varphi(x)$ satisfying
$\varphi^{\prime\prime}(x)+{2s{v}}(x)\varphi^{\prime}(x)-{m}(x)\varphi(x)=0,$ (3)
@@
where we absorbed $\eta$ by redefining the fields ${v}(x)\to{v}(x)/\eta$ and ${m}(x)\to{m}(x)/\eta$. We refer to the eigenvalue $s=\pm 1$ as the pseudospin. Notice that, if $\mathrm{Im}({m}(x))\neq 0$ or $\mathrm{Im}({v}(x))\neq 0$, the equation is not real and does not have real solutions in the general case. We recall that real wavefunctions are the hallmark of Majorana zero modes. In other words, the zero modes of a nonhermitian system are not necessarily Majorana zero modes.
@@ II.2 Line gap nonhermitian topology
Assuming uniform fields $v(x)=v$, $m(x)=m$, the Hamiltonian in Eq. 1 becomes diagonal in the momentum eigenfunctions and can be written as
@@
with an energy spectra given by $E(k)^{2}=(k^{2}+m)^{2}+4v^{2}k^{2}$. In the nonhermitian case, the energy spectra are generally complex. In nonhermitian systems shen_topological_2018 ; gong_topological_2018 ; kawabata_symmetry_2019 , energy spectra can exhibit either point gaps, where the energy eigenvalues do not cross a specific point in the complex plane, or line gaps, where the energy eigenvalues do not cross a specific line in the complex plane. nonhermitian spectra with line gaps can always be adiabatically deformed to hermitian and gapped spectra. The nonhermitian complex energy spectra of Eq. 4 shows a line gap, which is the line in the complex plane with $\mathrm{Im}(E(k))=0$, and can be deformed to hermitian real energy spectra with an energy gap by taking the limits $\mathrm{Im}(v)\to 0$, $\mathrm{Im}(m)\to 0$, as shown in Fig. 1. The line gap closes when $E(k)=0$ for some momenta $k\in\mathbb{R}$. Hence, the closing of the line gap is determined if the condition $E(k)=0$ is satisfied for any $k\in\mathbb{R}$, i.e., if any of the solutions of the algebraic equation $(k^{2}+m)^{2}+4v^{2}k^{2}=0$, which are given by $k=\pm\sqrt{-\left(\sqrt{m+v^{2}}\pm\sqrt{v^{2}}\right)^{2}}$, is real. This condition is satisfied only if $\mathrm{Re}\left(\sqrt{m+v^{2}}\pm\sqrt{v^{2}}\right)=0$. Hence, the line gap of the complex energy spectra closes when the quantity
@@
II.3 Exact solutions of zero modes at smooth domain walls
We assume that the fields become constant at large distances $x\to\pm\infty$ and that they approach their asymptotic values exponentially as $|{m}(x\to\mp\infty)-{m}_{L,R}|\sim{e}^{-|x|/w}$, $|{v}(x\to\mp\infty)-{v}_{L,R}|\sim{e}^{-|x|/w}$. Here, the constant $w>0$ is a characteristic length that measures the width of the smooth domain wall localized at the origin $x=0$. Under these very general assumptions, by mapping the whole real line $-\infty<x<\infty$ into the finite segment $10<y<$ 1 with the substitution $y(x)=(1+\tanh{(x/2w)})/2$, we found that for a given pseudospin $s=\pm 1$, the general solution of the modified Jackiw-Rebbi equation is given by a linear combination $A_{1}\varphi_{1}^{s}(x)+A_{2}\varphi_{2}^{s}(x)$ where

### v2

@@ Abstract
The bulk-boundary correspondence predicts the existence of boundary modes localized at the edges of topologically nontrivial systems. The wavefunctions of hermitian boundary modes can be obtained as the eigenmodes of a modified Jackiw-Rebbi equation. The bulk-boundary correspondence has also been extended to nonhermitian systems, which describe physical phenomena such as gain and loss in open and non-equilibrium systems. Nonhermitian energy spectra can be complex-valued and exhibit point gaps or line gaps in the complex plane, whether the gaps can be continuously deformed into points or lines, respectively. Specifically, line-gapped nonhermitian systems can be continuously deformed into hermitian gapped spectra. Here, we find the analytical form of the wavefunctions of nonhermitian boundary modes with zero energy localized at smooth domain boundaries between topologically distinct phases by solving the generalized Jackiw-Rebbi equation in the nonhermitian regime. Moreover, we unveil a universal relation between the scalar fields and the decay rate and oscillation wavelength of the boundary modes. This relation quantifies the bulk-boundary correspondence in nonhermitian line-gapped systems through physical quantities that are experimentally measurable. Furthermore, this relation is not affected by the specific spatial variations of the scalar fields. These results offer new insights into the localization properties of boundary modes in nonhermitian and topologically nontrivial states of matter.
@@ I Introduction
The bulk-boundary correspondence determines the presence and number of localized modes in relation to the topological invariants in topologically nontrivial condensed matter systems hatsugai_chern_1993 ; ryu_topological_2002 ; teo_topological_2010 , such as topological insulators and superconductors schnyder_classification_2009 ; hasan_colloquium:_2010 ; qi_topological_2011 ; shen_topological_2011 ; shen_topological_2017 . However, the bulk-boundary correspondence cannot predict the physical properties of the localized modes, such as their localization length, the presence or absence of oscillations in the wavefunction, and other local properties. Generalizations of the bulk-boundary correspondence apply to nonhermitian systems ashida_non-hermitian_2020 ; okuma_non-hermitian_2023 as well. Nonhermitian Hamiltonians naturally describe a variety of physical systems, such as open systems in the presence of gain and loss mechanisms, or dissipation, and exhibit the presence of complex energy spectra shen_topological_2018 ; gong_topological_2018 ; kawabata_symmetry_2019 ; ashida_non-hermitian_2020 ; okuma_non-hermitian_2023 ; schindler_hermitian_2023 ; nakamura_bulk-boundary_2024 in several different condensed matters and metamaterials gao_observation_2015 ; ghatak_observation_2020 ; zhang_observation_2021 ; su_direct_2021 ; fleckenstein_non-hermitian_2022 ; lin_observation_2022 ; liu_experimental_2023 ; wang_non-hermitian_2023 ; ochkan_non-hermitian_2024 . A fundamental distinction in this case is between spectra with line gaps and point gaps shen_topological_2018 ; gong_topological_2018 ; kawabata_symmetry_2019 . Point gaps describe energy eigenvalues that do not cross a reference point (usually the origin at zero). In contrast, line gaps describe energy eigenvalues that do not cross a reference line in the complex plane. Any nonhermitian Hamiltonian with line gaps can be adiabatically deformed into a corresponding hermitian Hamiltonian with conventional energy gaps. Hence, the bulk-boundary correspondence and the definition of topological invariants can be directly extended in this case. Hermitian Hamiltonians with nontrivial topology are naturally described by a modified Jackiw-Rebbi equation shen_topological_2011 ; shen_topological_2017 , a spinorial second-order linear differential equation. This generalizes the original Jackiw-Rebbi equation jackiw_solitons_1976 , which is instead of first-order. Henceforth, nonhermitian Hamiltonians can be described by a nonhermitian version of the modified Jackiw-Rebbi equation with complex coefficients.
@@
In recent works, we found the analytical wavefunctions of the zero energy modes of the modified Jackiw-Rebbi equation localized at smooth domain walls between topologically inequivalent phases in the hermitian regime marra_topological_2024 and extended some of these findings to the nonhermitian regime marra_zero_2025 . In this work, we derive the analytical form of the wavefunctions of zero modes in nonhermitian line-gapped systems described by a modified Jackiw-Rebbi equation with space-dependent mass and velocity terms, allowing for complex coefficients. The space dependency describes the case of a smooth domain wall, i.e., the case where the scalar fields are nonuniform within a region of finite length $w$, which we call a smooth domain wall, and become uniform at larger distances $x\gg w$. This allows us to derive the bulk-boundary correspondence for nonhermitian line-gapped systems by directly establishing the relation between nonhermitian topological invariants and the asymptotic decay rate of the zero modes. We thus analyze several properties of these zero modes, such as the localization length $\xi$, oscillation wavelength $\lambda$, conditions for the reality of the wavefunction, local properties, and local topological invariant, and show specific examples of nonhermitian zero modes corresponding to different regimes. This allows us to classify zero modes in different categories: Featureless zero modes in the limit $w\to 0$; Non-featureless zero modes with localization length and oscillation wavelength are larger than the width of the smooth domain wall $\xi,\lambda>w>0$; Non-featureless zero modes where the width of the smooth domain wall is larger than the localization length and oscillation wavelength $w>\xi,\lambda>0$. Featureless zero modes are described only by their localization lengths and oscillation wavelengths on the two sides of the smooth domain wall: For this reason, we call these zero modes with "no hair", since they can be completely described by only two numbers (in analogy with the terminology used for black holes). By contrast, we call non-featureless zero modes as zero modes with "hair". We can further distinguish zero modes with "short hair" for $\xi,\lambda>w>0$, which appear featureless at long length scales, and with "long hair" for $w>\xi,\lambda>0$, which appear non-featureless at all length scales. We also discuss possible experimental signatures of the bulk-boundary correspondence by unveiling a universal relation between the localization lengths, oscillation wavelengths, and bulk physical quantities.
II Exact wavefunctions of the zero energy modes
II.1 Modified Jackiw-Rebbi equation in the nonhermitian regime
@@
Figure 1: Energy spectra of the generalized Jackiw-Rebbi equation with uniform fields $v(x)=v$, $m(x)=m$ in Eq.˜3. (a) Hermitian case with $v,m\in\mathbb{R}$ corresponding to a real energy spectrum with an energy gap. (b) nonhermitian case with $v,m\notin\mathbb{R}$ corresponding to a complex energy spectrum with a line gap.
@@
The Jackiw-Rebbi equation jackiw_solitons_1976 is a Dirac equation where the mass term is space-dependent. Here, we consider a generalization of the Jackiw-Rebbi equation, where both the mass and Dirac velocity are space-dependent
@@
with $\Psi(x)$ a two-component spinor describing the fermionic field, $\sigma_{xyz}$ the Pauli matrices, $p=-\mathrm{i}\partial$ the momentum operator, and $\eta\geq 0$. Notice that the equation above differs from the original Jackiw-Rebbi equation by the addition of a quadratic term in the momentum. We consider here the nonhermitian case where the fields are complex-valued ${m}(x),{v}(x)\in\mathbb{C}$. This equation can be interchangeably referred to as a modified Jackiw-Rebbi equation or a modified Dirac equation with space-dependent mass and Dirac velocity. The hermitian case ${m}(x),{v}(x)\in\mathbb{R}$ describes topological insulators and superconductors and has been studied at length in our previous work Ref. marra_topological_2024 . The nonhermitian case $\mathrm{Im}({m}(x))\neq 0$ describes the presence of a nonhermitian potential V(x)=m(x)V(x)={m}(x). Examples of such systems are topological insulators perturbed by a space-dependent imaginary potential, describing gain and loss dynamics. In particular, smoothly varying, space-dependent gain and loss terms have been implemented in photonic lattices in several experiments wimmer_observation_2015 ; bandres_topological_2018 ; zhao_non-hermitian_2019 ; wang_hybridization_2025 . On the other hand, the nonhermitian case $\mathrm{Im}({v}(x))\neq 0$ describes, e.g., imbalanced superconducting pairing corresponding to dissipation effects on Cooper pairs. Examples of such systems are Kitaev chains with imbalanced superconducting pairing li_topological_2018 ; zhao_anomalous_2021 ; zhao_defective_2021 and other nonhermitian superconducting states ghatak_theory_2018 . Note that the term $\propto 2v(x)p\sigma_{y}$ does not describe the familiar case of a gauge-invariant hermitian $p$-wave superconducting pairing with arbitrary complex phase $\propto\mathrm{Re}(\Delta(x))p\sigma_{y}+\mathrm{Im}(\Delta(x))p\sigma_{x}=\Big(\begin{smallmatrix}0&-\Delta^{*}(x)\partial\\ \Delta(x)\partial&0\end{smallmatrix}\Big)$ but an imbalanced nonhermitian pairing $\propto\mathrm{Re}(\Delta(x))p\sigma_{y}+\mathrm{Im}(\Delta(x))p\sigma_{y}=\Big(\begin{smallmatrix}0&-\Delta(x)\partial\\ \Delta(x)\partial&0\end{smallmatrix}\Big)$ (we remind the reader that differentiation is an antihermitian operation). In the hermitian case, our model describes zero modes in one dimension with time-reversal, charge conjugation, and chiral symmetry, corresponding to the Altland-Zirnbauer class BDI (see Ref. marra_topological_2024 ). In the nonhermitian case, this corresponds to nonhermitian topological phases with real line gaps (see below) in one dimension in the Altland-Zirnbauer class BDI. However, since all nonhermitian topological phases can also be described in the continuum by a generalized Jackiw-Rebbi nonhermitian Hamiltonian shen_topological_2018 , these findings can in principle be extended to other nonhermitian topological phases with real line gaps, imaginary line gaps, and point gaps in all Altland-Zirnbauer classes in one or higher dimensions kawabata_symmetry_2019 .
We will consider the zero modes $E=0$ that are normalizable and satisfy the Dirichlet boundary conditions on infinite or semi-infinite intervals. Zero energy modes of Eq.˜1 are eigenstates of $\sigma_{x}$, i.e., $\sigma_{x}\Psi(x)=s\Psi(x)$, where the eigenvalue $s=\pm 1$ is referred to as the pseudospin. Hence, the zero modes are given by $\Psi(x)\propto\tiny\begin{pmatrix}1\\ s\end{pmatrix}\varphi(x)$, with $s=\pm 1$ and where $\varphi(x)$ satisfies
$\eta\varphi^{\prime\prime}(x)+{2s{v}}(x)\varphi^{\prime}(x)-{m}(x)\varphi(x)=0.$ (2)
@@
Hereafter we take $\eta=1$, without loss of generality, since the scalar fields can be always redefined as ${v}(x)\to{v}(x)/\eta$, ${m}(x)\to{m}(x)/\eta$. If $\mathrm{Im}({m}(x))\neq 0$ or $\mathrm{Im}({v}(x))\neq 0$, the equation is not real and does not have real solutions in the general case. We recall that real wavefunctions are the hallmark of Majorana zero modes. In other words, the zero modes of a nonhermitian system are not necessarily Majorana zero modes.
@@ II.2 Line gap nonhermitian topology
Assuming uniform fields $v(x)=v$, $m(x)=m$, the Hamiltonian in Eq.˜1 becomes diagonal in the momentum eigenfunctions and can be written as
@@
with an energy spectra given by $E(k)^{2}=(k^{2}+m)^{2}+4v^{2}k^{2}$. In the nonhermitian case, the energy spectra are generally complex. In nonhermitian systems shen_topological_2018 ; gong_topological_2018 ; kawabata_symmetry_2019 , energy spectra can exhibit either point gaps, where the energy eigenvalues do not cross a specific point in the complex plane, or line gaps, where the energy eigenvalues do not cross a specific line in the complex plane. Line gaps can be further classified as real line gaps, where the energy levels cannot cross the imaginary axis, and imaginary line gaps, where the energy levels cannot cross the real axis. Nonhermitian spectra with line gaps can always be adiabatically deformed to hermitian and gapped spectra. The nonhermitian complex energy spectra of Eq.˜3 shows a real line gap, which is the line in the complex plane with $\mathrm{Im}(E(k))=0$, and can be deformed to hermitian real energy spectra with an energy gap by taking the limits $\mathrm{Im}(v)\to 0$, $\mathrm{Im}(m)\to 0$, as shown in Fig.˜1. The line gap closes when $E(k)=0$ for some momenta $k\in\mathbb{R}$. Hence, the closing of the line gap is determined if the condition $E(k)=0$ is satisfied for any $k\in\mathbb{R}$, i.e., if any of the solutions of the algebraic equation $(k^{2}+m)^{2}+4v^{2}k^{2}=0$, which are given by $k=\pm\sqrt{-\left(\sqrt{m+v^{2}}\pm\sqrt{v^{2}}\right)^{2}}$, is real. This condition is satisfied only if $\mathrm{Re}\left(\sqrt{m+v^{2}}\pm\sqrt{v^{2}}\right)=0$. Hence, the line gap of the complex energy spectra closes when the quantity
@@
II.3 Exact wavefunctions
In this Section, we quickly derive the general solutions of the Jackiw-Rebbi equation in the nonhermitian regime by directly generalizing the solutions in the hermitian regime found in our previous work marra_zero_2025 . For a detailed derivation of these solutions, see Appendix˜A. We assume that the fields become constant at large distances $x\to\pm\infty$ and that they approach their asymptotic values exponentially as $|{m}(x\to\mp\infty)-{m}_{L,R}|\sim{e}^{-|x|/w}$, $|{v}(x\to\mp\infty)-{v}_{L,R}|\sim{e}^{-|x|/w}$. Here, the characteristic length $w>0$ measures the width of the smooth domain wall. Under these very general assumptions, using the substitution $y(x)=(1+\tanh{(x/2w)})/2$, we map the whole real line $(-\infty,\infty)$ into the finite segment $(0,1)$. This allows us to find the eigenmodes of the modified Jackiw-Rebbi equation given by $A_{1}\varphi_{1}^{s}(x)+A_{2}\varphi_{2}^{s}(x)$ with